STRUCTURAL, CONFORMATIONAL AND VIBRATIONAL STUDIES OF ISOCYANOCYCLOPENTANE FROM INFRARED, RAMAN SPECTRA AND AB INITIO CALCULATIONS

\underline{DATTATRAY K. SAWANT}, JOSHUA J. KlAASSEN, JAMES R. DURIG, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF MISSOURI-KANSAS CITY, MO 64110 USA.

The infrared and Raman spectra (3200 to 50 cm^{−1}) of the gas, liquid or solution, and solid have been recorded of isocyanocyclopentane, _{c}-C_{5}H_{9}NC. FT-microwave studies have also been carried out and 23 transitions were recorded for the envelope-axial (Ax) conformer. Variable temperature (-55 to -100^{∘}C) studies of the infrared spectra (3200 to 400 cm^{−1}) dissolved in liquid xenon have been carried out. From these data, both the Ax and envelope-equatorial (Eq) conformers have been identified and their relative stabilities obtained. The enthalpy difference has been determined to be 102 ±10cm^{−1} (1.21 ±0.03kJmol^{−1}) with the Ax conformer the more stable form. The percentage of the Eq conformer is estimated to be 38 ± 1% at ambient temperature. The conformational stabilities have been predicted from ab initio calculations by utilizing several different basis sets up to aug-cc-pVTZ from both MP2(full) and density functional theory calculations by the B3LYP method. Vibrational assignments have been made for the observed bands for both conformers with initial predictions by MP2(full)/6-31G(d) ab initio calculations to obtain harmonic force constants, wavenumbers, infrared intensities, Raman activities and depolarization ratios for both conformers. The heavy atom distances (Å): C≡N = 1.176 ; C-N≡C= 1.432; C-C_{β},C_{β}′ = 1.534; C_{β}-C_{γ}, C_{γ}′ = 1.542; C_{γ}-C_{γ}′ = 1.554 and angles (^{∘}):^{}C̸-N= 177.8; C̸_{β}C-N= 110.4; ^{}C̸_{β}CC_{β}′= 102.9; ^{}C̸C_{β}C_{γ} = 103.6; ^{}C̸_{β}C_{γ}C_{γ}′ = 105.9. The results are discussed and compared to the corresponding properties of some related molecules.